OpenStudy (anonymous):

The Room Is 4 Meters Longer Than It Is Wide. Its Perimeter Is 24 Meters. Find The Width In Meters

OpenStudy (anonymous):

What is the formula for the area?

OpenStudy (angela210793):

let l=length and w=width u have l=w+4 The perimeter of a rectangle is the sum of the four sides P=2(length)+2(width)=2(length+width) 24=2*(w+4+w) 24=2*(2w+4) 12=2w+4 12-4=2w 8=2w w=4

OpenStudy (anonymous):

2l+2w= P

OpenStudy (anonymous):

w=4

OpenStudy (anonymous):

Uh, perimeter is the sum of all the perimeter dimensions, both ends and both top and bottom. see drawing

OpenStudy (anonymous):

THANKS EVERYONE I FOUND MY ANSWER : )

OpenStudy (anonymous):

or in the case of a rectangle 2W + 2L = perimeter.